5. A college student borrows $1500 during her senior year. The loan is to be paid pack in equal monthly payments starting three years after the loan is established (interest accumulates during the grace period). The APR is 2.8% compounded monthly. How much should she expect to pay per month if she only wants to have to make payments for 2 years?

6. Mary has a credit card balance of $4,300 at an APR of 12% compounded monthly. Mary has elected to make interest only payments each month. (A) How much should she expect her monthly payments to be? (B) What will the balance be in her account after 2 years? (C) What would the balance be after 5 years? Mary realizes that this is a bad situation and decides that she wants to pay off the balance in one year’s time. (D) How much must she pay each month to have a zero balance in the account?
Answer #1
5 Amount borrowed $1,500
Monthly interest =(2.8/12)% 0.23333%
Monthly interest 0.002333
Future Value (FV)=Present Value (PV)*((1+i)^N)
Present Value=$1500
i=interest rate =0.0023333
N=number of months =36(3 years)
FV=Amount accumulated in three years $1,631.28 1500*(1.0023333^36)
Pv Amount at the end of 3 years $1,631.28
Rate Interest rate 0.23333%
Nper Number of months of payment 24 (2*12)
PMT Monthly payment required $69.97 (Using PMT function of excel with Rate=0.2333%, Nper=24, Pv=-1631.28)
6 Monthly interest =(12/12)% 1%
Credit Card Balance $4,300
(A)Monthly payment required=4300*1% $43
(B) balance in her account after 2 years $4,300
.(C)Balance in her account after 5 years $4,300
D
Pv Loan Balance $4,300
Rate Interest Rate(Monthly) 1%
Nper Number of months to clear the loan 12
PMT Amount she needs to pay each month $382.05 (Using PMT function of excel with Rate=1%, Nper=12, Pv=-4300)
016 =PMT(D14,D15,-D13) G H I J K L M D E F 5 Amount borrowed $1,500 Monthly interest =(2.8/12)% 0.23333% Monthly interest 0.0D28 : * fc =PMT(D26,027,-D25) A B D E F G H I J K L i=interest rate=0.0023333 N=number of months =36(3 years) FV=Amount accum